OpenStudy (anonymous):

Hydrogen sulfide gas is bubbled through excess potassium hydroxide solution. H2S + 2OH– → S2– + 2H2O How do you get there from H2S + KOH -> K2S + H2O ?

OpenStudy (chmvijay):

what exactly you want ???

OpenStudy (anonymous):

H2S + KOH -> K2S + H2O Write the balanced chemical equation excluding spectator ions. I want to know how to get from there to this: H2S + 2OH– → S2– + 2H2O

OpenStudy (chmvijay):

just separate the atoms from the molecules and give them charges example H2S can be written as 2 H+ + S^-2 like that u right for all molecules

OpenStudy (nincompoop):

here, you have a double replacement reaction. AB + CD = AD + CB first, the side of your reaction is incorrect. you were told that: Hydrogen sulfide gas is bubbled through excess potassium hydroxide solution, but you only put H2S + 2OH– → S2– + 2H2O notice that you are missing potassium on the reaction side. without the aid of a miracle, it would be difficult to balance if you do not account everything in the reaction.

OpenStudy (anonymous):

how do you get from here... H2S + KOH -> K2S + H2O to here... H2S + 2OH– → S2– + 2H2O

OpenStudy (nincompoop):

like I said it is a double replacement reaction AB + CD = AD + CB \[H_2 = A; S=B; K=C;(OH)=D\] \[[AD]=H_2+(OH)=H_2O\] \[[CB]= K+S = KS\] make sure it is balanced. it seems that you are doing net ionization problem and solubility. this involves what dissolves during a reaction and what will be solid (known as precipitate). like chmvijay indicated add their appropriate charges and use the solubility rules.